OpenStudy (anonymous):

Halogens have ------ electron affinity a. exothermic b. endothermic c . positive d . zero

OpenStudy (anonymous):

a

OpenStudy (kira_yamato):

There are 2 approaches to this. The first is to look at E˚ values for the halogens. The second would be to say that halogens form halides to attain the octet structure. Hence electron affinity is exothermic. \[E^0 = -\frac{∆G}{zF} = \frac{T∆S - ∆H}{zF}\]